# Paloma[Uncaptioned image] : A Benchmark for Evaluating Language Model Fit

### v1

@@ 2.4 Metrics
When comparing results where model vocabularies must differ, for instance research to improve tokenizers, Paloma follows Gao et al. (2020) in using bits per byte (BPB). This metric normalizes the log likelihood $\ell$ over documents by the count of UTF-8 encoded bytes in the corpus, $B$:
$\text{BPB}=\frac{1}{B}\text{log}_{2}(e^{-\ell})=\frac{-\ell}{B~{}\text{ln}(2)}$ (3)
@@
$\mu(\ell_{v})=\frac{1}{\mathbf{T}_{v}(N)}\sum_{t\in N}\sum_{i}^{\mid t\mid}\mathds{1}(v=t_{i})\,\text{ln}~{}p(t_{i}|t_{<i})$ (4)
2.4.1 Efficiency Metrics
In addition to performance metrics, we also ask submissions to Paloma to record measures of cost associated with the training of their language model: number of model parameters and number of tokens seen in training. We also record the size of the training dataset in UTF-8 encoded bytes and when models have run for more than one epoch—where increase in novel data ceases but training duration continues increasing. We elect to measure these abstract cost values rather than metrics of realized costs such as energy use or GPU hours, so that our efficiency comparisons are agnostic to hardware. Our aim is not to judge what hardware a submission uses. Note that this does not capture improvement from innovations that use hardware more efficiently. Such questions are better explored through benchmarks that control hardware such as Peng et al. (2023).
2.5 Comparable Submissions
Fair comparisons of language models can be challenging since there are so many variables to account for, like the number of parameters in each model, the amount of training data trained on, and the tokenizer used. In this section we highlight a number of ways that our benchmark can be used to provide evidence for practitioners to make scientific claims regarding how their model compares against other models.
We encourage submissions to Paloma to opt into the training guidelines in §2.1.1 (specifically as they are implemented in corresponding controls in §3.1). Likewise we request submissions not intending to study changes to vocabulary opt in to using the vocabulary of GPTNeoX-20B (Black et al., 2022). Where submissions opt out of these measures they will be marked for the corresponding limitations to comparability, allowing results to be filtered to remove all results that are not decontaminated, for instance.
Submissions can use inference code provided by us that supports any model integrated with Hugging Face Transformers (Wolf et al., 2020) to compute the metrics in §3 over each domain in the evaluation data. To make fair comparisons, it will be suggested that practitioners provide information on the measures of efficiency discussed in §2.4.1, such as model size. Similarly, submissions can record the name of the training dataset used. Finally, as model performance is typically best at the end of the learning rate schedule (compared with a model part way through training), the maximum duration of the learning rate schedule in tokens can be provided to mark comparability of partially trained checkpoints.
We outline a few types of scientific claims that can be made with our benchmark, including comparing different pretraining corpora and evaluating performance-efficiency tradeoffs:
  1. 1.
    When two models are trained on the same data, with the same cost budget (number of parameters or number of tokens seen), they can be directly compared. If one model outperforms the other, this is direct evidence that that model is better. This represents the most common type of comparison.
  2. 2.
    When two models have different computational budgets but achieve matching perplexities, this is evidence that the model with the lower computational cost is better. For example, if two models have matching perplexity, and are trained on the same number of tokens from the same corpus, where one model has fewer parameters, this is evidence that the smaller model is better.
  3. 3.
    When the model architecture, budget for number of training tokens, and other modeling configurations are fixed, and multiple training runs are done varying the training corpora,comparing the resulting trained models will effectively compare the pretraining corpora. This can provide evidence that one pretraining corpus is better than another. Our baseline experiments in this paper represent this type of scientific claim.
  4. 4.
    When a set of submissions fixes all configurations except for varying one dimension of cost (number of parameters or number of tokens seen), this can provide evidence of scaling trends for that model and training configuration.
3 Experimental Controls
In order to meet the guidelines we establish in §2.1, we implement a set of experimental controls whose technical details are discussed here. We further distinguish controls that must be applied during model training and controls that are applied at inference time.
3.1 Training Controls
3.1.1 Decontamination
@@
To mitigate contamination of our benchmark, we develop an approach for removing contamination from training data at the scale of pretraining corpora of trillions of tokens. We use a Bloom filter (Bloom, 1970) as implemented by Soldaini et al. (2023) to match training text that is contaminated with respect to the evaluation data. We employ this approach rather than the minHash or suffix array approaches used by Lee et al. (2022) and other deduplication work, as our approach is much more lightweight: the minHash approach would require pairwise computations, $O(|X_{t}||X_{e}|)$ between all training texts, $X_{t}$, and evaluation texts, $X_{e}$, where our approach runs a constant number of hashes, $K<<|\mathcal{X}_{e}|$, over all texts in $O\left(K(|X_{t}|+|X_{e}|)\right)$. Meanwhile the implementation of the suffix array approach of Lee et al. (2022) requires memory usage proportional to the size of the pretraining corpora. Since we aim to encourage researchers submitting to the benchmark to run this decontamination on their pretraining data, we opt to minimize cost and engineering complexity.
@@
Using our approach to find text matches, we mark contamination in the following way. We match text at the paragraph level, i.e., newline separated spans of text. This granularity strikes a balance between, on one hand, examining only full documents, which can miss contamination embedded in novel documents, and, on the other hand, all n-grams of a given size, where the size of the n-grams must be carefully set. Instead paragraph matching leverages this naturally occurring unit of language, although this heuristic has its own limitations especially in domains such as code or poetry, where line separation is handled very differently from prose. To avoid coincidental collisions in the space of small strings, we ignore matches in paragraphs smaller than 13 unicode segmented tokens (Unicode, 2023), as 13 is the n-gram sized used in contamination checks in Brown et al. (2020) and Rae et al. (2021). Similarly, we ignore paragraphs composed of only punctuation, spaces, and emoji, as, unlike words, these can be arbitrarily repeated when used as formatting, leading to high frequency n-grams greater than our 13-gram threshold. Lastly, as code data consists almost entirely of short and often repeated lines, we forgo any decontamination on these sources (Dolma-100-programming-languages and the The Stack domain of Dolma). We leave the question of how to properly decontaminate code data to future work.
@@
Having marked contaminated paragraphs, we now take the conservative measure of removing whole documents if they contain any contaminated paragraph. This has the added benefit of not disrupting the contiguity of text within documents, which excising paragraphs would do. Applying this approach to the datasets on which we train 6 baseline models results in the removal rates shown in Table 4. While these vary by orders of magnitude from dataset to dataset (with The Pile perhaps receiving a higher removal rate due to the intentional oversampling in that dataset), this approach removes at most $2.753\%$ of documents, making it feasible to apply without dramatically reducing training dataset size. Nevertheless, care should be taken to examine removal rates when applying this approach to new datasets.
3.1.2 Data Order
@@
As contemporary LMs train on instances that are themselves concatenations of training documents up to the maximum sequence length of the model, to fix the order of training data one cannot simply fix the order of documents but must train on the same concatenated instances. Achieving this requires not just a fixed random seed for training instance shuffling, but also adopting the same tokenization and maximum sequence length. Further fixing the number of instances in each gradient update would be required for fully identical training, however this is onerous for experiments that may be run on different hardware requiring different batch sizes. A compromise instead is to ensure that training code feeds instances into gradient steps in a deterministic shuffled order, so the relative ordering of data remains the same even if a given instance may fall in different gradient updates. In conclusion, we adopt the most direct way of controlling data order—we have submissions opting into this control use the same training code that we use to pretrain our baseline models. [^4] At the time of preprinting this training code will not yet be publicly released as it is developed under another project that has not yet concluded. Until it is released submissions wishing to opt in to this control should contact us for direct assistance with reproducing data order.
3.2 Evaluation Controls
3.2.1 Subsampling
@@
In Figure 2, we evaluate perplexity on data from C4 using Pythia 1.4B (Biderman et al., 2023) while varying the size of the evaluation subsample and training checkpoint. Each point in this figure represents the mean of perplexity on 20 different uniform subsamples and standard deviation is represented by the shaded region. As we expect, for a given checkpoint standard deviation shrinks as the evaluation subsample gets larger. More subtly, standard deviation shrinks as the model is trained on more data. This second observation matters if we want to measure model performance throughout training. Lastly note that the mean value is relatively stable over different evaluation subsample sizes, though a slight downward trend appears at the smallest subsample sizes.
@@
As a small first step towards more principled subsampling, we set the target subsample size based on the simplifying assumption that our metric variance results on C4 hold for other domains and models. Extrapolating our observations, we aim to subsample each split to a minimum of 1 million tokens per source and a minimum of 100 thousand tokens per domain. All datasets with domains are subsampled to 100 thousand tokens per domain other than Manosphere Corpus which we treat as a single domain, ICE which we include in entirety for comparability to its use in HELM, and Dolma which we subsample at a higher target of 500 thousand tokens per domain. A few sources fall below our thresholds, with Wikitext-103, Penn Treebank, and TwitterAAE being smaller than 1 million tokens per split despite being included in their entirety, and RedPajama having only 7 domains leading to 700 thousand tokens per split. We show the final token statistics in Table 3.
@@
If extrapolation from the trends we observed holds, perplexities on sources will be drawn from a distribution over subsamples with less than 1 standard deviation even at very early stages of training. Meanwhile, results on domains will be drawn for a similarly stable distribution by the end of training. This is admittedly a heuristic simplification, as the relationship between variability and subsampling will also likely depend on other factors such as average document length and heterogeneity of the source data, as well as the power of the model being evaluated. We must leave it to future benchmarks to explore these questions as the requirement of decontaminating pretraining data against evaluation data means any change to the evaluation data necessitates costly rerunning of pretraining of all baselines and submissions.
3.2.2 Vocabulary
@@
Where possible we control by the simplest approach of using the same vocabulary: the vocabulary used in GPT-NeoX-20B (Black et al., 2022) with 3 special tokens added by Dolma for masking personally identifiable information. Note that when vocabulary is fixed this is essentially a training control, as the model must be pretrained with this vocabulary. Nevertheless we mark this as an evaluation control, as we provide an option applied at inference time for making comparisons of models already pretrained with different vocabularies. Specifically, we follow The Pile (Gao et al., 2020) and use bits per byte (BPB; §3). In theory BPB may still present issues in comparability as it only includes likelihoods of the specific sequences produced by a given tokenizer, e.g., rain ##ing for the text raining, and not the marginal probability over all valid sequences in that vocabulary which would produce the identical text, e.g., ra ##in ##ing and so on (Mielke, 2019; Cao & Rimell, 2021; see also Hofmann et al., 2021). Models with a larger event space of possible sequences representing the same text will be at a disadvantage if they assign any non-zero probability to these valid predictions ignored by the metric. However, it has been shown empirically that the difference between the marginal probability over all valid sequences and the likelihood of the sequence produced by the tokenizer is small (Mielke & Eisner, 2018) and typically lower than 0.5% (Chirkova et al., 2023). So in conclusion, we encourage submissions to opt in to our fixed vocabulary and mark these as most comparable, but we also make allowance for submissions that opt out by only measuring comparisons involving models with different vocabularies in BPB.
3.2.3 Evaluation Format
@@
We follow the input format established by The Pile (Gao et al., 2020). In this format, documents are evaluated individually, e.g., “<BOS>document 1” then “<BOS>document 2”, rather than packed into concatenated maximum sequence length inputs, e.g., “<BOS>document 1<BOS>document 2<BOS>…”, where <BOS> is a special token for demarcating sequences. The latter concatenated approach is still often used as it takes the same preprocessing as is most commonly used for training data and is thus convenient for measuring validation loss during training. However, in Appendix §D we find preliminary evidence that the predictability of variance from subsampling observed in §3.2.1 breaks down for concatenated inputs. We also believe that evaluating documents individually more closely mirrors how models are used in practice at inference time. Providing more than one document at a time through concatenation is essentially a form of few shot in context learning for language modeling, as it allows the model to condition on information shared between concatenated documents when they are all drawn from the same domain. This is perhaps an interesting task formulation of its own but one that should be undertaken intentionally.
@@
Moreover, following The Pile, we split documents longer than maximum sequence length into disjoint inputs. This is also described by Press et al. (2021) as nonoverlapping inference. It is contrasted with sliding window inference in which some amount of overlapping tokens are included as context in maximum-sequence-length windows to prevent an unrealistic lack of conditioning for tokens in the middle of a document appearing shortly after a multiple of the maximum sequence length. However, a sliding window requires re-encoding overlapping tokens, making nonoverlapping inference the most efficient approach to computing perplexity.
4 Case Studies
@@
By applying our experimental controls (§3) to Paloma2), we are able to dig deeper into what language distributions models are learning to fit. In this section, we present several case studies demonstrating the types of analyses possible with Paloma. In §4.1, we use our 6 baseline 1B models that vary only in which common corpus they are pretrained on to isolate the effect of data composition on LM fit. In §4.2, we examine how scaling dynamics differ over the breadth of domains in Paloma. Finally in §4.3, we go beyond domains and decompose perplexity by performance on different vocabulary types (i.e., specific elements of the model vocabulary).
We hypothesize that one of the strongest drivers of differences in performance between different domains is the composition of the pretraining data of a language model. While we show in §4.2 that scaling model parameters or tokens seen increases performance on nearly all domains, the pretraining data composition directly determines the distribution of language that the model is learning to fit, which may or may not align with the distributions of language in the domains we evaluate. Therefore we examine the impact of varying the pretraining corpus while holding all other experimental decisions the same.
Figure 3: Ordinary perplexity over all the standard language modeling and fine-grained domain sources in Paloma, excluding code data not supported in our decontamination. While this aggregation obscures subtler differences in performance between domains, for uses such as monitoring training stability, this metric provides a single measure of fit over more diverse data than is typically used for measuring validation loss (e.g., C4).
Ordinary perplexity
In Figure 3, we consider the most simple and aggregated view of LM fit that Paloma can provide—an ordinary perplexity as defined in §3. Specifically we compute perplexity over all data in the standard language modeling and fine-grained domain sources. The other sources are set aside for now as they are designed for targeted analysis of questions such as the fit of models to discourse with prevalent toxicity. We also exclude the code data in Dolma and Dolma-100-programming-languages, which is not supported by our decontamination approach. Using this view we can already see that the baseline models which are trained only on data derived from Common Crawl (C4, Falcon Refinedweb, and mC4-en) stand out from the other baselines which also incorporate more curated sources of data. However, this also points to the limitation of this most aggregated view of the results: this ordinary perplexity represents fit to domains in proportion to the number of tokens we have chosen to sample from each domain. As we sample 100 thousand tokens from each domain and the majority of our domains are not sourced from Common Crawl, that data source is much less represented in Paloma than most of the pretraining corpora whose held-out data is conventionally used to measure validation loss. Nevertheless this simplified view of the results is useful for specific use cases that need a single metric over a prescriptive mix that emphasizes robustness to a diversity of domains, largely derived from non-web scraped sources.
Figure 4: For each source with domains, domain perplexity for the final checkpoint of each model ordered by median domain perplexity over all models. While performance gaps between some baselines are highly consistent across domains (e.g., RedPajama and The Pile baselines on Dolma-100-subreddits), others exhibit noisy performance gaps per domain that do not follow the trend in median domain difficulty (e.g., the mC4-en baseline on C4-100-domains). Note that these erratic-gap patterns are frequently on the baselines pretrained on just Common Crawl data.
Macro average perplexity
In Figure 1, we provide another aggregation that examines the robustness of fit by considering all domains equally—a macro average of perplexity over domains: $|D|^{-1}\sum_{d\in D}\text{perplexity}(d)$ for domains, $D$. By contrast the previous ordinary perplexity is essentially an exponentiated micro average over the domains implicitly selected for during corpus curation. Macro averaging lets all marked domains have equal say on the model’s performance, instead. To make these macro averages more easily interpretable, we can examine them separately per source. The most striking pattern that emerges here is the high, and sometimes non-monotonic, perplexity of the 3 baselines trained on only Common Crawl data (C4, mC4-en, Falcon Refinedweb). One source where this is most apparent is evaluating on The Pile. There the Falcon Refinedweb and mC4-en baselines’ results are dominated by greater than 10,000 perplexity on the Ubuntu IRC domain, while other domains are in the low tens, and the C4 baseline exhibits an identical pattern but with 8,056 perplexity on ArXiv. Both these domains contain large amounts of non-natural language, in the form of LaTeX and shell code as well as angle-bracketed IRC usernames. So while these Common Crawl baselines spike on different domains, it appears they are all more susceptible to these extreme gaps in fit to some domains, perhaps due to a lack of exposure to non-natural language such as code or otherwise due to having only one set of cleaning filters applied to a single source of data.
Perplexity per domain ordered by median perplexity
Rather than aggregating, we can visualize each domain perplexity separately to surface gaps in fine-grained LM fit. In Figure 4, we arrange the domains by their median perplexity over the baselines, as this order gives some sense of the intrinsic difficulty of a domain. We can then see which baselines more or less follow this order, differing only by a consistent offset, and which have gaps that are more idiosyncratic to each domain. Again we see that when baselines have irregular gaps from the median these are most frequently baselines pretrained on only Common Crawl. The notable exception is The Pile baseline on M2D2 S2ORC and Dolma-100-programming-languages, which has erratic gaps substantially below the median, perhaps indicating that baseline is benefiting from exposure to specific domains and not others rather than only a overall facility for scientific papers and code. The erratic-gapped Common Crawl baselines, by contrast, are all worse than median perplexity, suggesting that they may have complete gaps in exposure to features of certain domains that cannot be recovered through generalization.
4.2 Scaling Improves Domain Fit Unequally

### v2

@@ Appendix B Additional Metrics
When comparing results where model vocabularies must differ, for instance research to improve tokenizers, Paloma follows Gao et al. [2020] in using bits per byte (BPB). This metric normalizes the log likelihood $\ell$ over documents by the count of UTF-8 encoded bytes in the corpus, $B$:
$\text{BPB}=\frac{1}{B}\text{log}_{2}(e^{-\ell})=\frac{-\ell}{B~{}\text{ln}(2)}$
@@
$\mu(\ell_{v})=\frac{1}{\mathbf{T}_{v}(N)}\sum_{t\in N}\sum_{i}^{\mid t\mid}\mathds{1}(v=t_{i})\,\text{ln}~{}p(t_{i}|t_{<i})$
Appendix C Experimental Controls
Here we discuss the details of the experimental controls (introduced in §3) that we implement to meet our guidelines for rigorous perplexity evaluations. We distinguish controls that must be applied during model training and controls that are applied at inference time.
C.1 Training Controls
C.1.1 Decontamination
@@
A basic tenet of machine learning is that for evaluation to accurately represent performance, training and test data need to be non-overlapping. However, large pretraining corpora are known to contain evaluation data and large models are known to memorize training data [Dodge et al., 2021, Elazar et al., 2023, Carlini et al., 2022]. Lee et al. [2022] show in their second figure that models underestimate perplexity on evaluation documents with near duplicates in the training corpus by several points relative to models with those duplicate training documents removed. Thus benchmarks of language modeling should actively remove contaminated training data, rather than just partitioning held out splits by documents, assuming no documents overlap. The Pile applies document-level deduplication to two of their 22 domains before splitting held-out data, but its designers note that this does not prevent leakage of evaluation data more generally [Gao et al., 2020]. Furthermore, spans of contaminated text within larger unrelated documents can still contribute to overestimation of performance, so decontamination should be conducted at a sub-document level. To our knowledge, Paloma is the first language modeling benchmark to require removing training data that is contaminated with respect to evaluation data.
@@
To mitigate contamination of our benchmark, we develop an approach for removing contamination from training data at the scale of pretraining corpora of trillions of tokens. We use a Bloom filter [Bloom, 1970] as implemented in Soldaini et al. [2024] to match training text that is contaminated with respect to the evaluation data. We employ this approach rather than the minHash or suffix array approaches used by Lee et al. [2022] and other deduplication work, as our approach is much more lightweight: the minHash approach would require pairwise computations, $O(|X_{t}||X_{e}|)$ between all training texts, $X_{t}$, and evaluation texts, $X_{e}$, where our approach runs a constant number of hashes, $K<<|\mathcal{X}_{e}|$, over all texts in $O\left(K(|X_{t}|+|X_{e}|)\right)$. Meanwhile the implementation of the suffix array approach of Lee et al. [2022] requires memory usage proportional to the size of the pretraining corpora. Since we aim to encourage researchers using our benchmark to run this decontamination on their pretraining data, we opt to minimize cost and engineering complexity.
@@
Using our approach to find text matches, we mark contamination in the following way. We match text at the paragraph level, i.e., newline separated spans of text. This granularity strikes a balance between, on one hand, examining only full documents, which can miss contamination embedded in novel documents, and, on the other hand, all n-grams of a given size, where the size of the n-grams must be carefully set. Instead paragraph matching leverages this naturally occurring unit of language, although this heuristic has its own limitations especially in domains such as code or poetry, where line separation is handled very differently from prose. To avoid coincidental collisions in the space of small strings, we ignore matches in paragraphs smaller than 13 unicode segmented tokens [Unicode, 2023], as 13 is the n-gram sized used in contamination checks in Brown et al. [2020] and Rae et al. [2021]. Similarly, we ignore paragraphs composed of only punctuation, spaces, and emoji, as, unlike words, these can be arbitrarily repeated when used as formatting, leading to high frequency n-grams greater than our 13-gram threshold. Lastly, as code data consists almost entirely of short and often repeated lines, we forgo any decontamination on these sources (Dolma-100-programming-languages and the The Stack domain of Dolma). We leave the question of how to properly decontaminate code data to future work.
@@
Having marked contaminated paragraphs, we now take the conservative measure of removing whole documents if they contain any contaminated paragraph. This has the added benefit of not disrupting the continuity of text within documents, which excising paragraphs would do. Applying this approach to the datasets on which we train 6 baseline models results in the removal rates shown in Table 4. While these vary by orders of magnitude from dataset to dataset (with The Pile perhaps receiving a higher removal rate due to the intentional oversampling in that dataset), this approach removes at most $2.753\%$ of documents, making it feasible to apply without dramatically reducing training dataset size. Nevertheless, care should be taken to examine removal rates when applying this approach to new datasets.
Another limitation arises from our use of documents as a fundamental unit of data. This impacts our decontamination approach, since we remove whole documents that have any paragraph marked as contaminated to avoid mangling documents by excising individual paragraphs. Such an approach tends to disproportionately remove long documents that are frequently quoted, which may include seminal works (e.g., Martin Luther King’s “I Have a Dream” speech) that actually deployed models should be familiar with. The purpose of Paloma, however, is to enable controlled research on the science of language modeling, but production models should likely use caution in applying this decontamination technique.
C.1.2 Data Order
@@
As contemporary LMs train on instances that are themselves concatenations of training documents up to the maximum sequence length of the model, to fix the order of training data one cannot simply fix the order of documents but must train on the same concatenated instances. Achieving this requires not just a fixed random seed for training instance shuffling, but also adopting the same tokenization and maximum sequence length. Further fixing the number of instances in each gradient update would be required for fully identical training, however this is onerous for experiments that may be run on different hardware requiring different batch sizes. A compromise instead is to ensure that training code feeds instances into gradient steps in a deterministic shuffled order, so the relative ordering of data remains the same even if a given instance may fall in different gradient updates. In conclusion, we adopt the most direct way of controlling data order—we recommend using the same training code that we use to pretrain our baseline models.
C.2 Evaluation Controls
C.2.1 Subsampling
@@
In Figure 5, we evaluate perplexity on data from C4 using Pythia 1.4B [Biderman et al., 2023] while varying the size of the evaluation subsample and training checkpoint. Each point in this figure represents the mean of perplexity on 20 different uniform subsamples and standard deviation is represented by the shaded region. As we expect, for a given checkpoint standard deviation shrinks as the evaluation subsample gets larger. More subtly, standard deviation shrinks as the model is trained on more data. This second observation matters if we want to measure model performance throughout training. Lastly note that the mean value is relatively stable over different evaluation subsample sizes, though a slight downward trend appears at the smallest subsample sizes.
@@
As a small first step towards more principled subsampling, we set the target subsample size based on the simplifying assumption that our metric variance results on C4 hold for other domains and models. Extrapolating our observations, we aim to subsample each split to a minimum of 1 million tokens per source and a minimum of 100 thousand tokens per domain. All datasets with domains are subsampled to 100 thousand tokens per domain other than Manosphere Corpus which we treat as a single-domain source, ICE which was included in early versions of Paloma in entirety for comparability to its use in HELM, and Dolma which we subsample at a higher target of 500 thousand tokens per domain. A few sources fall below our thresholds, with WikiText-103, Penn Treebank, and TwitterAAE being smaller than 1 million tokens per split despite being included in their entirety, and RedPajama having only 7 domains leading to 700 thousand tokens per split. We show the final token statistics in Table 1.
@@
If extrapolation from the trends we observed holds, perplexities on sources will be drawn from a distribution over subsamples with less than 1 standard deviation even at very early stages of training. Meanwhile, results on domains will be drawn for a similarly stable distribution by the end of training. This is admittedly a heuristic simplification, as the relationship between variability and subsampling will also likely depend on other factors such as average document length and heterogeneity of the source data, as well as the power of the model being evaluated. We must leave it to future benchmarks to explore these questions as the requirement of decontaminating pretraining data against evaluation data means any change to the evaluation data necessitates costly rerunning of pretraining of all baselines.
Another limitation arises from our use of documents as a fundamental unit of data. When subsampling although we balance the number of tokens used to represent each domain, we still sample documents until that target token count is reached. Concretely, this means that some domains, especially books, are represented by only dozens of documents, which likely does not capture the full distribution of the domain as well as many smaller documents might.
C.2.2 Vocabulary
Perplexity per token is not comparable between models with different vocabularies [Jelinek, 1998] or, by extension, different tokenizers [Mielke, 2019]. Since models distribute probability over a vocabulary of tokens, models with larger vocabularies will tend to have higher perplexities than ones with smaller vocabularies. Where possible, the most rigorous solution is to impose one vocabulary on all experiments, allowing perplexity to be directly compared. Some lines of research, such as improving tokenizers, require comparisons of LM fit across vocabularies. This is possible by normalizing likelihood by a segmentation intrinsic to the text such as characters or bytes [Mielke, 2019]. The Pile [Gao et al., 2020] proposes BPB (Appendix B) as the best compromise when tokenizers are not identical, an approach we adopt as well. Paloma further establishes a standard tokenizer and vocabulary for experiments that do not need to change this experimental variable.
@@ C.2.2 Vocabulary
Where possible we control by the simplest approach of using the same vocabulary: the vocabulary used in GPT-NeoX-20B [Black et al., 2022] with 3 special tokens added by Dolma for masking personally identifiable information. Note that when vocabulary is fixed this is essentially a training control, as the model must be pretrained with this vocabulary. Nevertheless we mark this as an evaluation control, as we provide an option applied at inference time for making comparisons of models already pretrained with different vocabularies. Specifically, we follow The Pile [Gao et al., 2020] and use BPB. In theory BPB may still present issues in comparability as it only includes likelihoods of the specific sequences produced by a given tokenizer, e.g., rain ##ing for the text raining, and not the marginal probability over all valid sequences in that vocabulary which would produce the identical text, e.g., ra ##in ##ing and so on (Mielke, 2019, Cao and Rimell, 2021; see also Hofmann et al., 2021). Models with a larger event space of possible sequences representing the same text will be at a disadvantage if they assign any non-zero probability to these valid predictions ignored by the metric. However, it has been shown empirically that the difference between the marginal probability over all valid sequences and the likelihood of the sequence produced by the tokenizer is small [Mielke and Eisner, 2018] and typically lower than 0.5% [Chirkova et al., 2023]. So in conclusion, we encourage those using Paloma to opt in to our fixed vocabulary, or make comparisons involving models with different vocabularies in BPB.
C.2.3 Evaluation Format
While perplexity is clearly defined as a function of the likelihood assigned by a model to a set of sequences, the manner in which that likelihood is computed may vary depending on how inputs are formatted for the model. The Pile [Gao et al., 2020] identify one possible variation: inferring test documents as separate inputs or concatenating them together to fill a single input. Meanwhile, Press et al. [2021] point out that documents larger than the maximum sequence length can be split either with or without overlap.
@@ C.2.3 Evaluation Format
We follow the input format established by The Pile [Gao et al., 2020]. In this format, documents are evaluated individually, e.g., “<BOS>document 1” then “<BOS>document 2”, rather than packed into concatenated maximum sequence length inputs, e.g., “<BOS>document 1<BOS>document 2<BOS>…”, where <BOS> is a special token for demarcating sequences. The latter concatenated approach is still often used as it takes the same preprocessing as is most commonly used for training data and is thus convenient for measuring validation loss during training. However, in Appendix H we find preliminary evidence that the predictability of variance from subsampling observed in Appendix C.2.1 breaks down for concatenated inputs. We also believe that evaluating documents individually more closely mirrors how models are used in practice at inference time. Providing more than one document at a time through concatenation is essentially a form of few shot in context learning for language modeling, as it allows the model to condition on information shared between concatenated documents when they are all drawn from the same domain. This is perhaps an interesting task formulation of its own but one that should be undertaken intentionally.
@@
Moreover, following The Pile, we split documents longer than maximum sequence length into disjoint inputs. This is also described by Press et al. [2021] as nonoverlapping inference. It is contrasted with sliding window inference in which some amount of overlapping tokens are included as context in maximum-sequence-length windows to prevent an unrealistic lack of conditioning for tokens in the middle of a document appearing shortly after a multiple of the maximum sequence length. However, a sliding window requires re-encoding overlapping tokens, making nonoverlapping inference the most efficient approach to computing perplexity.
Appendix D Additional Case Studies
@@
In this section, we present additional case studies to explore analyses possible with Paloma. Previously in §4.1, we use our 6 baseline 1B models that vary only in which common corpus they are pretrained on to isolate the effect of data composition on LM fit. In §4.2 we introduced the observation that most loss occurs on the most common vocabulary types, which we now expand on in Appendix D.2 by analyzing performance dynamics of different vocabulary types. First, in Appendix D.1, we examine how scaling dynamics differ over the breadth of domains in Paloma.
Results in the appendix include two additional sources The Pile [Gao et al., 2020] and ICE [Greenbaum and Nelson, 1996], however access restrictions on these datasets prevent us from rehosting them. As such we have removed them from the body of our paper, but still share our findings on these datasets as auxiliary results not part of Paloma.
D.1 Scaling Improves Domain Fit Unequally